OpenStudy (anonymous):

Please check my work to see if O am correct on minimum yvallue on the graph of y=f(x) f(x)=-x^2+6x+8 6/-2=3 x=f(3) (3)^2+6*(3)+8 3^2=9 6*3=18 +8 y-value is 35?

OpenStudy (anonymous):

6/-2=3 I don't know what you did here. (3)^2+6*(3)+8 You are missing the negative sign out in front of x^2 here

OpenStudy (anonymous):

6/-2=3 The answer to this is -3 but still don't know why you did this

OpenStudy (anonymous):

I went and found the x coordinate of the vertex so b=6 a=-2

OpenStudy (anonymous):

because it is -x^2 there is only going to be a maximum. No minimum

OpenStudy (anonymous):

because you would 6/-2*1

OpenStudy (anonymous):

\[f(x)=-x^2+6x+8\]\[f'(x)=-2x+6\]\[0=-2x+6\]\[-6=-2x\]\[x=3\]\[f(3)=-(3)^2+6(3)+8\]\[f(3)=17\]\[\max=(3,17)\]

OpenStudy (anonymous):

Are you allowed to use Calculus? it is so much easier